A ship travels between two ports. The cost of fuel is 100 (ax+bx+10), when x is the average speed of the ship in knots, and a and b are constants. If the ship travels at 4 knots, the cost of fuel would be 9000, but at 6 knots, the cost would be 7000.
a) find the values of a and b
b) what is the cost of fuel at an average speed of 5 knots?

Answer by Alan3354(69443)

A ship travels between two ports. The cost of fuel is 100 (ax+bx+10), when x is the average speed of the ship in knots, and a and b are constants. If the ship travels at 4 knots, the cost of fuel would be 9000, but at 6 knots, the cost would be 7000.
a) find the values of a and b
b) what is the cost of fuel at an average speed of 5 knots?
==============
C = 100x*(a+b) + 1000
9000 = 400(a+b) + 1000
a+b = 20
-----------------
C = 100x*(a+b) + 1000
7000 = 600(a+b) + 1000
a+b = 100
------------------
a & b are not constants.
No solution.